Solve
d^2y/dx^2 + 1/x×dy/dx= 12Inx/x^2
1
Expert's answer
2020-11-03T17:35:04-0500

Let us solve the equation "\frac{d^2y}{dx^2}+\frac{1}{x}\frac{dy}{dx}=12\frac{\ln x}{x^2}".


Let us multiply both part by "x": "x\frac{d^2y}{dx^2}+\frac{dy}{dx}=12\frac{\ln x}{x}". Then


"\frac{d}{dx}(x\frac{dy}{dx})=12\frac{\ln x}{x}" and therefore,


"x\frac{dy}{dx}=12\int \frac{\ln x}{x}dx=12\int \ln x\ d(\ln x)=6\ln^2x+C_1"


Let us divide both part by "x":


"\frac{dy}{dx}=6\frac{\ln^2x}{x}+\frac{C_1}{x}"


Consequently,


"y=6\int(\frac{\ln^2x}{x}+\frac{C_1}{x})dx=6\int\ln^2x\ d(\ln x)+6C_1\int\frac{dx}{x}=2\ln^3 x+6C_1\ln x+C_2"


Answer: "y=2\ln^3 x+6C_1\ln x+C_2"